It takes an older computer 4 times as long to send out a company’s email as it does a newer computer. Working together, it takes the two computers
10 minutes to send out the email. How long will it take the newer computer to send out the email on its own?

Answer by stanbon(75887)

It takes an older computer 4 times as long to send out a company’s email as it does a newer computer. Working together, it takes the two computers
10 minutes to send out the email. How long will it take the newer computer to send out the email on its own?
-------------------------
Together rate = 1/10 job/min
---
Older computer rate = 1/(4x) job/min
----
Newer computer rate = 1/x job/min
---------------------------
Equation:
rate + rate = together rate
1/x + 1/4x = 1/10
Multiply thru by 20x to get:
20 + 5 = 2x
----
2x = 25
x = 12.5 minutes (time for the newer computer to do the job)
---
4x = 50 minutes (time for the older computer to do the job)
----------
Cheers,
Stan H.
------------------